# LLM-assisted Explicit and Implicit Multi-interest Learning Framework for Sequential Recommendation

### v1

@@ 4.2.2. Performance experiment on different backbones
To explore whether the EIMF framework is compatible, we designed the following experiment. We selected three classic models as backbone, which include the traditional single-interest SR models SASRec and Bert4Rec, and the multi-interest model MIND. We compared the recommendation performance of the above models with the model combined with the EIMF framework on the Beauty and Office datasets. The experimental results are shown in the Table4.
As can be seen from Table 4, the performance of all backbone models has been improved after adding the EIMF framework, which proves the wide applicability of the EIMF framework. Our analysis suggests that this is because EIMF greatly enriches the information resources on which behavioral interest modeling relies by introducing semantic interests, effectively compensating for the limitations due to data sparsity in traditional modeling approaches. In addition, EIMF adopts an inference mechanism based on typical samples, which not only greatly reduces the amount of data for LLM reasoning tasks but also further explores and demonstrates the positive impact of group interests on individuals.
(a) Office dataset
(b) Grocery dataset
(c) Beauty dataset
@@
(a) Office dataset
(b) Grocery dataset
(c) Beauty dataset
@@
4.3. Ablation Study
To verify the effectiveness of framework components, we designed two variants, EIMF w/o Align and EIMF w/o Predict, which represent the deleted modality alignment task and deleted semantic prediction task respectively. We experimented with the two variants and the whole framework on three datasets, and the results are shown in Figure 3.
As can be seen from Figure 3, the whole EIMF is the best-performing model on all three datasets, and the other two variants have varying degrees of performance degradation, which proves that the two auxiliary tasks we designed can effectively enhance the user’s multi-interest representation. In addition, the performance of the other two variants drops to different degrees on different datasets. In the Office and Beauty datasets, the performance of EIMF w/o Align drops the most, while in the Grocery dataset, the performance of EIMF w/o Predict drops the most. We believe that this is because different categories of items correspond to different user behavioral patterns, e.g., when purchasing office supplies or beauty products, users’ interests tend to change frequently; while when purchasing daily groceries, users’ interests are more purposive. Therefore, in the Grocery dataset, the invisible interests reflected by the behavioral patterns are more homogeneous, and the semantic prediction auxiliary task dominated by explicit interests can identify more potential interests, on the contrary, in the Beauty and Office datasets, the behavioral patterns have been able to reflect the diversified interests of the users, and therefore it is more necessary to enhance the invisible user behavioral interests by using the modal alignment task. The two auxiliary tasks complement each other and enable EIMF to adapt to more scenarios.
@@ 4.4. Hyperparameter Study
We conducted hyperparameter experiments on three datasets to explore the impact of preference degree ($p$) on the framework recommendation performance in the AP clustering algorithm. The $p$ can control the number of clusters in the AP algorithm. The smaller the $p$, the fewer clusters there are. The value of $p$ is set to $\{None,-2,-3,-4,-5,-7,-10\}$, where $None$ means not using a clustering algorithm to infer the interests of each user individually. In addition, corresponding to $p=\{-2,-5,-7,-10\}$, the typical number of samples obtained from clustering about the proportion of the total number of data are $\{80\%,50\%,20\%,5\%\}$, respectively.
@@
It can be seen from Figure 4 that the best p-values are different on different datasets. Specifically, on the Office dataset, when $p=-2$, the $\text{Recall}@50$ of EIMF(REMI) is 0.2492, and the $\text{NDCG}@50$ is 0.1010; on the Grocery dataset, when $p=-5$, the $\text{Recall}@50$ is 0.2856, and the $\text{NDCG}@50$ is 0.1214; on the Beauty dataset, when $p=-10$, the $\text{Recall}@50$ is 0.3636, and the $\text{NDCG}@50$ is 0.1348. Through analysis and discussion, we believe that the optimal p-value selection is closely related to the amount of data and data distribution. Specifically, when the amount of data is small or the behavioral similarity between users is weak, a larger p-value helps to identify as many typical samples as possible and ensure the independence and uniqueness of user interests; when the amount of data is large or the behavioral similarity between users is high, a smaller p-value can not only effectively reduce the computational burden faced by LLM when processing large amounts of data, but also help to discover a wider range of group characteristics and enrich user interest modeling.
@@ 5. Conclusions and Future Work
In this paper, we answer the question of how to combine large&\&&small models in RS from the perspective of multi-interest learning and semantic alignment. We propose an LLM-assisted framework with explicit and implicit multi-interest learning to capture users’ interests from both behavioral and semantic levels. Specifically, the framework consists of an implicit behavioral interest module and an explicit semantic interest module. The implicit behavioral interest module adopts the traditional multi-interest RS model and learns user behavioral interest patterns by analyzing the interaction data between users and items; the explicit semantic interest module, in the training phase, first classifies users based on the AP algorithm to select typical samples of each category, and then uses LLM to infer the typical samples to obtain multiple semantic interests. These semantic interests are subsequently used to augment the representation of user behavioral interests through auxiliary tasks, including semantic prediction and modality alignment. Observed from extensive experiments on real-world datasets, our EIMF framework can effectively and efficiently combine LLM with RS models, significantly improving the recommendation performance. This two-layer interest modeling method not only considers the interest preferences directly expressed by users but also deeply explores their potential tendencies, which can provide users with more comprehensive and accurate personalized recommendations. In future research, we plan to combine multimodal and multi-behavioral data further to explore the more fine-grained user interests reflected in these data.

### v3

@@ 5.2.3. Performance Experiment on Different Backbones
To explore whether the EIMF framework is compatible, we designed the following experiment. We selected three classic models as the backbone, which include the traditional single-interest SR model - BERT4Rec, and the multi-interest models MIND and REMI. We compared the recommendation performance of the above models with the model combined with the EIMF framework on the Beauty, Office, and Ml-1M datasets. The experimental results are shown in the Table 5.
As can be seen from Table 5, the performance of all backbone models has been improved after adding the EIMF framework across the Beauty, Office, and Ml-1M datasets, demonstrating the wide applicability of the EIMF framework. Specifically, on the Beauty dataset, EIMF enhances BERT4Rec’s Recall@20 from 0.1405 to 0.1421 and NDCG@20 from 0.078 to 0.0804, with similar improvements for MIND (Recall@20: 0.1563 to 0.1755) and REMI (Recall@20: 0.2189 to 0.2322). On the Office dataset, EIMF boosts BERT4Rec’s Recall@20 from 0.1037 to 0.1186 and NDCG@20 from 0.0635 to 0.0719, with MIND and REMI showing comparable gains. For the Ml-1M dataset, which features diverse movie preferences, EIMF significantly improves BERT4Rec’s Recall@20 from 0.1294 to 0.1405 and NDCG@20 from 0.3643 to 0.4011, while MIND improves from 0.0605 to 0.0654 (Recall@20) and REMI from 0.0834 to 0.0897 (Recall@20). Our analysis suggests that this consistent performance uplift is due to EIMF’s ability to enrich behavioral interest modeling by integrating semantic interests inferred by the LLM, effectively compensating for data sparsity and noise in traditional approaches. This effect is particularly pronounced in the Ml-1M dataset, where the diversity of user preferences (e.g., across movie genres) benefits from the semantic augmentation provided by ESIM. Additionally, EIMF’s inference mechanism, based on typical samples, reduces the computational burden of LLM inference while leveraging group-level interest patterns to enhance individual representations, further contributing to its effectiveness across all datasets and backbones.
5.3. Ablation Study
To validate the effectiveness of the framework components, we designed three variants: EIMF w/o Align, EIMF w/o Predict, and EIMF w/o Tar, which correspond to the removal of the modality alignment task, the semantic prediction task, and the target-aware attention mechanism, respectively. These variants were compared against the complete EIMF framework on three datasets—Office, Grocery, and Beauty—using metrics such as Recall@$20$ and NDCG@$20$. The results, illustrated in Figure 4, provide insights into the contributions of each component.
As can be seen from Figure 4, the complete EIMF framework consistently outperformed all variants across the three datasets, confirming the synergistic contributions of the modality alignment task, semantic prediction task, and target-aware attention mechanism. Below, we analyze the impact of each component:
  • EIMF w/o Align: Removing the modality alignment task caused the most significant performance drop on the Office and Beauty datasets. For instance, Recall@$20$ on Beauty decreased by approximately 7% compared to the whole model. We analyze that this is because the modality alignment task ensures consistency between implicit behavioral interest representations and explicit semantic interest representations, which is critical for datasets with diverse user interests, such as the Office Supplies dataset (e.g., different office supply needs) and the Beauty dataset (e.g., covering both makeup and skincare). Without alignment, the model struggles to integrate semantic signals into behavioral representations, leading to suboptimal recommendations. This effect is less pronounced in Grocery, where user interests are more homogeneous, suggesting that alignment is particularly vital for complex interest patterns.
  • EIMF w/o Predict: The removal of the semantic prediction task led to the most significant performance drop on the Grocery dataset, with Recall@$20$ decreasing by approximately 6%, while the impact on the Office and Beauty datasets was relatively minor. We believe this is because the semantic prediction task leverages semantic interests inferred by the LLM to uncover latent user preferences, which is particularly effective in the Grocery dataset, where users tend to exhibit purposeful behaviors (e.g., consistently purchasing daily necessities). By using semantic signals to predict the next item, this task enhances the model’s ability to capture subtle interests that are not fully reflected in behavioral data. In contrast, the Office and Beauty datasets feature inherently diverse behavior patterns, so the absence of this task has a smaller impact, as behavior modeling alone can partially compensate for the lack of semantic guidance.
  • EIMF w/o Tar: Removing the target-aware attention mechanism leads to a slight performance drop across all datasets. For example, Recall@$20$ decreases by approximately 4% and 3.5% on the Beauty and Office datasets, respectively, while the drop is smaller on the Grocery dataset, around 2.5%. The target-aware attention mechanism dynamically selects the most relevant interest representations with respect to the target item in both IBIM (using item ID embeddings) and ESIM (using textual embeddings). In the Beauty dataset, where user interests are highly diverse (e.g., eye makeup vs. skincare), this mechanism ensures accurate recommendations by focusing on contextually relevant interests, preventing unrelated items from dominating the output. In the sparser Office dataset, it helps reduce noise by prioritizing meaningful behavioral signals. In the more stable-interest Grocery dataset, the mechanism is less critical but still beneficial for optimizing interest selection and improving precision. Removing it results in broader and less targeted interest representations, increasing the likelihood of recommending irrelevant items.
(a) Office
(b) Grocery
(c) Beauty
@@ 5.3. Ablation Study
(a) Office
(b) Grocery
(c) Beauty
@@
(a) Office
(b) Grocery
(c) Beauty
Figure 6. The results of alignment task with different ($\alpha$, $\beta$) hyperparameter settings on three datasets.
(a) Office
(b) Grocery
(c) Beauty
Figure 7. The results of joint learning with different $\gamma$ hyperparameter settings on three datasets.
@@ 5.4. Hyperparameter Study
We conducted sensitivity experiments on the hyperparameters ($p$, $(\alpha,\beta)$, $\gamma$) involved in the framework on three datasets to explore their impact on the recommendation performance. The following is an introduction to the experimental settings for each hyperparameter:
@@
  • The $p$ can control the number of clusters in the AP algorithm. The smaller the $p$, the fewer clusters there are. The value of $p$ is set to $\{None,-2,-3,-4,-5,-7,-10\}$, where $None$ means not using a clustering algorithm to infer the interests of each user individually. In addition, corresponding to $p=\{-2,-5,-7,-10\}$, the typical number of samples obtained from clustering about the proportion of the total number of data are $\{80\%,50\%,20\%,5\%\}$, respectively.
  • ($\alpha,\beta$) can control the influence ratio of the two loss functions in the modality alignment task, and ($\alpha,\beta$) is set to { (0.1, 0.4), (0.2, 0.3), (0.3, 0.2), (0.4, 0.1) }.
  • $\gamma$ can control the influence of the auxiliary task in joint learning, $\gamma$ is set to $\{0.01,0.03,0.05,0.1,0.2\}$.
5.4.1. The Impact of Clustering Preference $p$
@@
It can be seen from Figure 5 that the best p-values are different on different datasets. Specifically, on the Office dataset, when $p=-2$, the $\text{Recall}@50$ of EIMF is 0.2492, and the $\text{NDCG}@50$ is 0.1010; on the Grocery dataset, when $p=-5$, the $\text{Recall}@50$ is 0.2856, and the $\text{NDCG}@50$ is 0.1214; on the Beauty dataset, when $p=-10$, the $\text{Recall}@50$ is 0.3636, and the $\text{NDCG}@50$ is 0.1348. We believe that when the amount of data is small or the behavioral similarity between users is weak, a larger p-value helps to identify as many typical samples as possible and ensure the independence and uniqueness of user interests; when the amount of data is large or the behavioral similarity between users is high, a smaller p-value can not only effectively reduce the computational burden faced by LLM when processing large amounts of data, but also help to discover a wider range of group characteristics and enrich user interest modeling.
5.4.2. The Impact of $\alpha$ and $\beta$ in the modality alignment Task
Figure 6 shows the performance of the different ($\alpha,\beta$) settings on three datasets, from which we can see that the ratio of $\alpha$ to $\beta$ has different effects on different datasets, and obvious peaks can be found on each sub-figure. In the Office and Beauty datasets, when $\alpha$ and $\beta$ are set to 0.4 and 0.1, the framework achieves the best performance, with $\text{Recall}@50$ metrics of 0.2296 and 0.3636 respectively. For the Grocery dataset, the best performance was achieved with $\alpha$ set to 0.1 and $\beta$ set to 0.4, with $\text{Recall}@50$ metric of 0.2964.
The variation in optimal $(\alpha,\beta)$ settings highlights the interplay between contrastive loss and cosine similarity loss in aligning modalities, tailored to dataset-specific behavioral patterns: The Office and Beauty datasets exhibit diverse and complex user interests (e.g., office supplies range from stationery to electronics, and beauty products span cosmetics and skincare). A higher $\alpha=0.4$ prioritizes the contrastive loss, which excels at distinguishing positive pairs (where behavioral and semantic representations align) from negative ones, thereby enhancing the model’s ability to capture subtle interest differences. A lower $\beta=0.1$ ensures that the cosine similarity loss serves as a supplementary constraint—strengthening alignment without over-smoothing the representations, which could obscure fine-grained interest distinctions. The dominance of contrastive loss aligns with the need for robust separation between different interests, as it encourages representations to be discriminative rather than merely similar. A higher $\beta$ would overemphasize cosine similarity, potentially leading to overly generalized representations that fail to capture the diversity of user interests, thereby reducing recommendation accuracy.
In contrast, user behavior in the Grocery dataset is more purposeful and homogeneous (e.g., regular purchases of daily necessities). Here, a higher $\beta=0.4$ emphasizes the cosine similarity loss, which minimizes the angular distance between behavioral and semantic representations, creating a unified embedding space well-suited to consistent interest patterns. A lower $\alpha=0.1$ reduces the influence of contrastive loss, as the need to distinguish between diverse interests is less pronounced. A higher $\alpha$ would overweight the contrastive loss, which could force the model to over-differentiate representations, introducing noise in a dataset with relatively uniform interests.
5.4.3. The Impact of $\gamma$ in the Joint Learning
Figure 7 depicts the effect of different $\gamma$ settings on performance. According to the figure, in the three datasets examined, the framework’s performance consistently reaches the optimal level when $\gamma$ is equal to 0.1. This phenomenon shows that setting $\gamma$ to 0.1 is the best choice for the degree of influence of the auxiliary task: it can ensure that it does not interfere with the main task, and can assist the main task by enhancing representation learning, thereby improving the recommendation accuracy.
When $\gamma$ is too low (e.g., 0.01), the model tends to behave like a traditional sequential recommender, relying heavily on behavioral data and ignoring the rich semantic knowledge provided by the LLM. This limits its ability to handle diverse or sparse interest patterns. Conversely, when $\gamma$ is too high (e.g., 0.2), the auxiliary tasks dominate training, which may introduce noise from semantic prediction or lead to overfitting on aligned representations—effects that do not necessarily translate into better recommendation performance.
5.5. Case Study
Figure 8. An example of interest reasoning using user click record (left) and typical sample click record (right).
(a) Direct Inference Interests
(b) Typical Samples Interests
Figure 9. Word clouds comparing interests inferred directly from user clicks (a) and from typical samples (b).
In this section, we explore the critical role of typical samples within the framework through a case study. Taking the user “ID 12” from the Beauty dataset as an example, Figure 8 shows the user’s original click history on the left, consisting of 6 items, and the right side shows the click history of the typical sample corresponding to the user, which includes 43 items. The red-highlighted area indicates items in the typical sample that are similar to those clicked by the user “ID 12”, while the blue-highlighted area represents items the user “ID 12” has not interacted with but may potentially be interested in.
In Figure 9, we use word clouds to visualize the results of inference based on the user’s own historical records and those based on the historical records of typical samples. In comparison, we can see that the typical samples contain a significantly wider range of interests, including product types that users have not yet explicitly expressed but may be interested in (although users have not interacted, but users with similar interests have interacted with the products). For example, users who are interested in eye makeup may also be in demand for eyelash care products. Such association analysis helps to reveal the user’s potential needs and thus enriches the user’s interest representation. The AP clustering algorithm can help us find typical samples that are highly similar to other samples in a specific category. Such typical samples not only represent the core features of the category but also because their data volume is usually the richest in the category, which means that they can capture a wider range of user preferences and behavior patterns.
@@ 6. Conclusions and Future Work
In this paper, we propose a bi-level multi-interest learning framework that effectively exploits the strengths of LLM and SR models. The implicit behavioral interest module adopts the traditional SR model and learns user behavioral interest patterns by analyzing the interaction data between users and items; the explicit semantic interest module, in the training phase, first classifies users based on the AP algorithm to select typical samples of each category, and then uses LLM to infer the typical samples to obtain multiple semantic interests. These semantic interests are subsequently used to augment the representation of user behavioral interests through auxiliary tasks, including semantic prediction and modality alignment. As observed from extensive experiments on real-world datasets, our EIMF framework can effectively and efficiently combine LLM with RS models, significantly improving the recommendation performance. This bi-level interest modeling method not only considers the interest preferences directly expressed by users but also deeply explores their potential tendencies, which can provide users with more comprehensive and accurate personalized recommendations.
In the future, we plan to explore combining multiple behaviors (purchase, collection, etc.) to reveal deeper user interests and combine techniques such as Retrieval Augmentation Generation (RAG) and Chain of Thoughts (CoT) to improve LLM’s reasoning ability about interests.
Acknowledgements.
The Australian Research Council supports this work under the streams of Future Fellowship (Grant No. FT210100624), the Discovery Project (Grant No. DP240101108), and the Linkage Project (Grant No. LP230200892).